OpenStudy (bacon):

The Length of one side of a rectangle is twice as long as the radius of a circle. You want the area of the rectangle to equal the area of the circle. How would I solve?

OpenStudy (anonymous):

The other side should be equal to circumference.

OpenStudy (bacon):

Why is it (2r) in the 3rd equation and not just r?

OpenStudy (bacon):

@Justanotherguy

OpenStudy (phi):

I would translate Length of one side of a rectangle is twice as long as the radius to L (length of one side) = (is) 2 * (twice as long) r (the radius) or L= 2r the area of the rectangle is W*L = W*2*r area of the circle is pi r^2 so we have 2rW= pi r^2

OpenStudy (phi):

to solve for W in \[ 2rW= \pi r^2\] divide both sides by 2r \[ \frac{2r W}{2r} = \frac{\pi r^2}{2r} \] and simplify

OpenStudy (anonymous):

@Bacon Sorry, I had to leave for some work. And thanks @phi !